OpenStudy (anonymous):

solve for 0 x^2-x+1=0

OpenStudy (cwrw238):

i think you mean x?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

would it be 0?

OpenStudy (cwrw238):

use the quadratic formula x = [-(-1) +/- sqrt(1 - 4 *1 *1) ] / 2 there are no real roots - they are complex

hero (hero):

Solve for 0...lol

OpenStudy (anonymous):

\[x^3+1=0\] \[(x+1)(x^2-x+1)=0\] \[x=-1, x=\frac{1}{2}\pm\frac{\sqrt{3}}{2}i\]

OpenStudy (cwrw238):

x = 1/2 + i sqrt3/2 and 1/2 - i sqrt3/2

OpenStudy (anonymous):

i dont understand any of it

OpenStudy (cwrw238):

i'm not sure where satellites coming from ? but we agree on the 2 comp;ex roots

OpenStudy (cwrw238):

haven't you come across complex numbers yet nikkiem?

OpenStudy (cwrw238):

if not then you should'nt be given this problem - maybe you typed it incorrectly?

OpenStudy (anonymous):

yeah but i never really understand them to much sorry im not go at math at all

OpenStudy (cwrw238):

well - all squares of real numbers are positive 2*2 = 4 , -3 * -3 = 9 etc A negative number h as no real square root so someone came along and suggested that theres am imaginative number sqrt(-1) which we cal 'i'. so square root of, say, sqrt( -9 ) = sqrt(9)* sqrt(-1) = = 3i

OpenStudy (anonymous):

so my answer then would be -1 right

OpenStudy (cwrw238):

no the imaginary number in the answer is sqrt (1 - 4) = sqrt (-3) = sqrt3 * sqrt(-1) = sqrt3 * i or as i wrote it i sqrt3 - i put the i first because it is sqrt 3 not sqrt (3i) the full answers are 1/2 + i sqrt3 / 2 and 1/2 - i sqrt3 / 2 its best to write them as satellite did using the equation thing but im not good at that.